# Wild2Avatar: Rendering Humans Behind Occlusions

### v1

@@ 3.3 Optimization Objectives
where $\alpha_{occ},\alpha_{fg}\in[0,1]$ are the obstacle and human density maps respectively, $\epsilon=0.1$ is a threshold to determine the occupancy of the human, $\texttt{BCE}(\cdot)$ is the binary cross entropy function, and $\texttt{sg}(\cdot)$ indicates gradient stopping. The workflow is outlined in Figure 4. We noticed that occlusions only appear in a small area of most videos, and the remaining areas usually have empty obstacle densities. To deal with this data imbalance, we use weighted $\texttt{BCE}(\cdot)$ and assign a higher weight to the pixels that tend to be occluded.
@@
where $\mathcal{X}_{near}$ can be determined by the Euclidean distances to the human mesh, which was initialized as SMPL [31] mesh and updated in the optimization process [9].
Our Wild2Avatar is trained on the combination of all of the above objectives, each with a scale weight $\lambda_{(\cdot)}$.
@@ 4 Experiments
Figure 5: Qualitative comparisons with Vid2Avatar [9] in the OcMotion dataset [13] (top 5) and in-the-wild videos (bottom 2). We manually blurred the faces for the in-the-wild videos to maintain anonymity.
@@ 4.1 Datasets
OcMotion [13]. This dataset consists of indoor scenes of humans engaging with a variety of objects while being partially occluded by them. We used 5 out of the 48 videos at different levels of occlusion to evaluate the methods. In particular, we drew only 100 frames from each of the videos to train the models. We initialized the optimization process with the dataset’s provided camera matrices, human poses, and SMPL parameters. Frame-wise binary human segmentation masks are obtained from the Segment Anything Model (SAM) [23].
In-the-wild videos. We conducted additional experiments on two real-world videos, one of which was downloaded from YouTube while the other one was captured by our research team using a cell phone camera. We sampled  150 frames from both of the videos for training. In these videos, we obtained camera matrices, human poses, and SMPL parameters using SLAHMR [7, 50]. Since no ground truth poses are provided, evaluations on these videos also indicate the robustness of methods to inaccurately estimated priors.
OcMotion $\texttt{quality}_{vis}$ comp. $\texttt{quality}_{llm}$
Vid2Avatar [9] 13.14 0.77 4.40
Wild2Avatar 13.07 0.81 8.20
In-the-wild $\texttt{quality}_{vis}$ comp. $\texttt{quality}_{llm}$
Vid2Avatar [9] 8.20 0.64 4.0
Wild2Avatar 9.45 0.73 7.5
Table 1: Quantitative comparisons on the 5 OcMotion and 2 in-the-wild videos. We color cells that have the best metric values.
@@ 4.2 Evaluations
Comparison. Wild2Avatar is mainly compared against Vid2Avatar [9], which is the state-of-the-art in human rendering. We also compare against the most recent occlusion-aware human rendering method OccNeRF [46]. For fairness, all comparison methods also use the same binary human segmentation masks either for supervision or ray sampling. Note that OccNeRF originally requires more than 500 frames for stable training while Wild2Avatar and Vid2Avatar only require $\sim$100. We conducted experiments on both high and low frame count regimes for OccNeRF.
@@
Metrics. We used qualitative and quantitative evaluations to compare the methods. For qualitative evaluations, we render the human behind the obstacles as well as novel views to assess the capability of occlusion handling. Furthermore, we rely on three quantitative metrics to measure the capability of renderers from different perspectives. First, we compute the commonly used Peak Signal-to-Noise Ratio (PSNR) on the visible human parts (determined by the binary segmentation mask) to measure the rendering quality on unoccluded human appearance. Then, we calculate the Intersection over Union (IoU) between human occupancy masks and GT SMPL mesh silhouettes to measure the completeness of rendering [46]. Lastly, we propose to assess the complete rendering quality of occluded humans through Large Language Models (LLMs). We prompt GPT4V [37] to output a quality score at a scale of 0-10 (the higher the better) for each rendering. The three metrics are denoted as $\texttt{quality}_{vis}$, comp., and $\texttt{quality}_{llm}$ respectively.
4.3 Results on Occluded Monocular Videos
@@
We first compare the rendering results of Vid2Avatar and Wild2Avatar on both datasets in Figure 5. Both methods can reconstruct the visible human parts in high fidelity without obvious artifacts or floaters, which verified the superiority of representing humans via implicit SDF. However, Vid2Avatar fails to recover the occluded body parts and mistakenly renders parts of the occlusions as human appearance. Wild2Avatar, on the other hand, not only decouples most occlusions from the human body but recovers consistent human appearance. In Table 1, we report the quantitative results of the two methods and observed on-par rendering performances at the visible part. Wild2Avatar surpasses Vid2Avatar consistently on body geometry and the rendering quality of the occluded parts.
4.4 Comparison Against OccNeRF [46]
@@
We then compare Wild2Avatar against a recent occlusion-aware human rendering counterpart OccNeRF [46] in Figure 6. For fair comparisons, we trained OccNeRF on 500 frames and 100 frames respectively. Without implicit SDF representation, OccNeRF suffers from common defects of floaters and artifacts. Although OccNeRF is also able to recover occluded human parts, the body is usually unexpectedly twisted resulting in low rendering quality.
OcMotion $\texttt{quality}_{vis}$ comp. $\texttt{quality}_{llm}$
OccNeRF-100 [46] 13.77 0.73 3.50
OccNeRF-500 [46] 15.02 0.77 5.00
Wild2Avatar 15.92 0.82 8.50
Figure 6: Comparison against OccNeRF [46] trained on 100 and 500 frames. We color cells that have the best metric values.
@@ 4.5 Visualizations of Scene Decomposition
4.6 Ablation Studies
Figure 8: Ablation results. Major differences are highlighted.
Impact of occlusion-aware parameterization. One of the major contributions that enables Wild2Avatar to achieve human renderings under occlusions is the occlusion-aware scene parameterization. Here we first ablate the effectiveness of such a design by inheriting the two-stage parameterization but optimizing the rendering of the foreground human on the visible parts only [46] while maintaining the same optimization objectives. As shown in the first row of Figure 8, even though Wild2Avatar can still recover the occluded appearances without the proposed parameterization, the rendered results are corrupted with many artifacts.
Impact of $\mathcal{L}_{occ}$. In section 3.3, the occlusion decoupling loss is proposed to further disentangle the rendering of human and occlusions. We validated the effect of this loss by simply removing it from optimization objectives. According to the results in the second row of Figure 8, the occluded areas cannot be fully recovered without the proposed loss.
Impact of $\mathcal{L}_{comp}$. As also noted in [46], the completeness of 3D human geometry needs to be explicitly enforced during training. When removing this loss, the 2D renderings can easily degenerate. We also observed that the proposed loss serves as a regularizer to force the human geometry to be consistent with the SMPL mesh prior, which prevents from rendering incorrect poses (third row in Figure 8).
@@ 5 Discussion and Conclusion
Discussion. Training human neural radiance fields to fit input frames while handling occlusions is challenging. Without knowing any semantic priors of the occlusions, it is impractical to achieve occlusion-human decoupling in a fully unsupervised manner. Our proposed occlusion decoupling loss $\mathcal{L}_{occ}$ serves as a weakly-supervised approach with the help of precomputed binary human segmentation masks. Moreover, we noticed that both Vid2Avatar and Wild2Avatar rely on accurate poses. Although the poses can be jointly optimized along the training process, inaccurate priors are still very likely to lead to poor or incomplete renderings, especially at body limbs. Moreover, the need to render the occlusions causes the inference time of Wild2Avatar to increase, resulting in a slower optimization process. Future work may include adding additional pose correction steps [50] and adopting efficient NeRF variants [18, 34, 22] for faster training.
@@ Acknowledgment.
This work was partially funded by the Panasonic Holdings Corporation, the Gordon and Betty Moore Foundation, the Jaswa Innovator Award, Stanford HAI, and Stanford Wu Tsai Human Performance Alliance.

### v2

@@ 3.3 Optimization Objectives
Figure 5: Qualitative comparisons with Vid2Avatar [1] on 6 occluded sequences in the OcMotion dataset [57].
@@
where pixels in $\alpha_{occ},\alpha_{fg}\in[0,1]$ are the obstacle and human densities respectively, $\epsilon=0.1$ is a threshold to determine the occupancy of the human, $\texttt{BCE}(\cdot)$ is the binary cross entropy function, and $\texttt{sg}(\cdot)$ indicates gradient stopping. The workflow is outlined in Figure 4. We noticed that occlusions only appear in a small area of most videos, and the remaining areas usually have empty obstacle densities. To deal with this data imbalance, we use weighted $\texttt{BCE}(\cdot)$ and assign a higher weight to the pixels that tend to be occluded.
@@
where $\mathcal{X}_{near}$ can be determined by the Euclidean distances to the human mesh, which was initialized as SMPL [29] mesh and updated during optimization [1].
Wild2Avatar is trained to optimize the combination of all of the above objectives, each with a scale weight $\lambda_{(\cdot)}$:
$\mathcal{L}_{rgb}+\lambda_{eik}\mathcal{L}_{eik}+\lambda_{dec}\mathcal{L}_{dec}+\lambda_{occ}\mathcal{L}_{occ}+\lambda_{comp}\mathcal{L}_{comp}.$ (8)
@@ 4 Experiments
Figure 6: Comparisons with Vid2Avatar [1] on 3 occluded in-the-wild monocular videos. Occluded humans from novel views are not available. We manually blurred the faces for the in-the-wild videos to maintain anonymity.
Figure 7: Comparisons with HSR [40] on 2 synthesized occluded videos. Occluded humans from novel views are not available.
@@ 4.1 Datasets
Unlike our prior work that was majorly evaluated on simulated occlusions [8], we focus more on evaluations on real-world occlusions in this extension work.
OcMotion [57]. This dataset consists of indoor scenes of humans engaging with a variety of objects while being partially occluded by them. We used 6 challenging sequences out of the 48 videos at different levels of occlusion to evaluate the methods. In particular, we drew only 100 frames from each of the videos to train the models. We initialized the optimization process with the dataset’s provided camera matrices, human poses, and SMPL parameters. Frame-wise binary human segmentation masks are obtained from the Segment Anything Model [58].
In-the-wild videos. We conducted additional experiments on 3 real-world videos, two of which was downloaded from YouTube with CC license while the other one was captured by our research team using a cell phone camera. We sampled  150 frames from all of the videos for training. In these videos, we obtained camera matrices, human poses, and SMPL parameters using SLAHMR [59, 60]. Since no ground truth poses are provided, evaluations on these videos also indicate the robustness of methods to inaccurately estimated priors.
Synthesized videos. To better evaluate our model under challenging conditions such as moving cameras and long-horizon occlusions, we synthesized occluded variants of 5 of the NeuMan [6] videos by inserting textured, box-shaped static occluders into each scene, guided by the 3D human trajectory and camera parameters from COLMAP [61]. These occluders are 3D-consistent with the background and rendered directly into the frames. Additional experiments were conducted on the 5 synthesized videos, bringing the total to 14—representing the most extensive evaluation of occluded human modeling with moving cameras to date. Similar to in-the-wild videos, we obtained SMPL parameters using SLAHMR [60].
Due to the scarcity of occluded human videos captured by moving cameras, most of our evaluations were conducted on videos recorded with stationary cameras, except for one in-the-wild video, whose result is shown in the bottom row of Figure 6. However, as an extension of Vid2Avatar [1], our method is not limited to fixed-camera scenarios and can be naturally applied to moving-camera setups.
4.2 Implementation Details
Following the common NeRF protocol, we optimize human NeRFs on a per-video basis. Our $\mathcal{F}^{fg}$ consists of two sub-networks: one for learning the implicit SDF (implicit network) and one for learning the point-wise radiance (rendering network). The implicit network uses a MLP with 8 layers of 256 feature dimensions and the rendering network uses a MLP with 4 layers of 256 feature dimensions. Similar to $\mathcal{F}^{fg}$, $\mathcal{F}^{scene}$ also has two sub-networks but in different configurations: the implicit network uses a MLP with 8 layers of 256 feature dimensions and the rendering network uses a single-layer MLP of 256 feature dimension. Both background and occlusion learn per-frame latent codes of 32 feature dimensions. We use only 64 ray samples for training and 128 ray samples for evaluation/inference [1]. During training, 512 rays are randomly sampled within the 2D human bounding box. Note that we use the same binary human segmentation mask for both Vid2Avatar and Wild2Avatar for fairness. The loss weights are set as $\lambda_{eik}=0.1$, $\lambda_{dec}=0.003$, $\lambda_{comp.}=0.2$, and $\lambda_{occ}=0.1$. We adopted the Adam [62] optimizer with an initial learning rate of $5.0e^{-4}$, which is decayed by $\frac{1}{2}$ at the 200th and the 500th step.
OcMotion In-the-wild
Methods $\texttt{quality}_{vis}$ comp. $\texttt{quality}_{user}$ $\texttt{quality}_{vis}$ comp. $\texttt{quality}_{user}$
[…]
TABLE I: Quantitative comparisons on the three datasets. We color cells that have the best metric values.
[…]
HSR [40] 18.77 0.9005 68.80
Wild2Avatar 22.47 0.9624 44.36
TABLE II: Quantitative comparisons on the synthesized videos. We color cells that have the best metric values.
@@ 4.3 Evaluations
Comparisons. Wild2Avatar is mainly compared against Vid2Avatar [1], which is the state-of-the-art in human rendering. We also compare against the previous occlusion-aware human modeling counterpart - OccNeRF [8]. For fairness, all comparison methods use the same binary human segmentation masks either for supervision or ray sampling. Note that OccNeRF originally requires more than 500 frames for stable training while Wild2Avatar and Vid2Avatar only require $\sim$100. We conducted experiments on both high and low frame count regimes for OccNeRF in Figure 8.
NeuMan [6], HOSNeRF [48], and HSR [40] are similar methods that also use layered representations of the scene. However, they have significant limitations preventing us from fully utilizing them as baselines. NeuMan learns a scene NeRF model, which is not feasible when the camera is stationary. On the other hand, the official HOSNeRF implementation cannot extend to custom videos beyond the six videos presented in their paper, and fails when the occlusion is not connected to the human. Additionally, both NeuMan and HOSNeRF require more resources than our method: NeuMan requires 3 GPUs for a total of 175 hours of training time, HOSNeRF requires 4 GPUs for a total of 118 hours of training time, while Wild2Avatar requires only a single 24GB GPU for 48 hours of training time. To better evaluate Wild2Avatar, we provide comparison results against HSR on 2 of the synthesized videos, which are shown in Figure 7 and quantified in Table II.
@@
Metrics. We used qualitative and quantitative evaluations to compare the methods. For qualitative evaluations, we render the human behind the obstacles as well as novel views to assess the capability of occlusion handling. Furthermore, we rely on three quantitative metrics to measure the capability of renderers from different perspectives. First, we compute the commonly used Peak Signal-to-Noise Ratio (PSNR) on the visible human parts (determined by the binary segmentation mask) to measure the rendering quality on unoccluded human appearance. Then, we calculate the Intersection over Union (IoU) between human occupancy masks and GT SMPL mesh silhouettes to measure the completeness of rendering [8]. Lastly, for a more rational assessment of the rendering quality of occluded humans, we conducted user studies to ask 25 out-of-field people to rate human renderings at a scale of 0-10 (the higher the better). The three metrics are denoted as $\texttt{quality}_{vis}$, comp., and $\texttt{quality}_{user}$ respectively.
For evaluations on the synthesized videos, we follow the common protocol and calculated PSNR and SSIM values between humans in the unoccluded raw videos and the rendered humans at the input views.
Figure 8: Comparison against OccNeRF [8] trained on 100 and 500 training frames.
Figure 9: Geometry reconstructions. We manually blurred the faces for the in-the-wild videos for maintaining anonymity.
4.4 Results
@@
We first compare the results of Vid2Avatar and Wild2Avatar on the OcMotion videos in Figure 5, where both methods demonstrate high fidelity in reconstructing the visible human parts without obvious artifacts or floaters, showcasing the superiority of representing humans via implicit SDF. However, Vid2Avatar struggles with recovering the occluded body parts and erroneously renders parts of the occlusions as human appearance. Wild2Avatar effectively decouples occlusions from the human body and recovers consistent human appearances. Figure 6 shows results on more challenging in-the-wild videos. Wild2Avatar consistently outperforms Vid2Avatar in real-world scenarios.
@@
Subsequently, we extend our comparison to include Wild2Avatar against the recent occlusion-aware human modeling method OccNeRF [8], as illustrated in Figure 8. For a fair comparison, OccNeRF was trained on datasets of 500 frames and 100 frames, respectively. Despite its ability to recover occluded human parts, OccNeRF, is prone to common defects such as floaters and artifacts, often resulting in the body appearing unexpectedly twisted and thus leading to a lower rendering quality.
In Table I, the quantitative results reveal on-par rendering performances for visible parts, with Wild2Avatar consistently outperforming Vid2Avatar in terms of body geometry and the rendering quality of occluded parts.
Moreover, we include comparisons between Wild2Avatar and HSR [40] - the most recent method that also decouples occlusions from human subjects - on the synthesized videos. As shown in Figure 7 and Table II, our rendering results exhibit significantly higher fidelity, with fewer occlusion artifacts and more complete human bodies. These results confirm that Wild2Avatar not only performs well on videos with moving cameras but also outperforms the current state-of-the-art.
4.5 Geometry Reconstructions
When rendering the human avatars through volumetric rendering, we can not only obtain their appearances but also reconstruct their underlying geometries. We overlay the reconstructed human surface normals on the input frames in Figure 9. The results show that Wild2Avatar excels in reconstructing human geometries from occluded videos.
4.6 Ablation Studies
Figure 11: Ablation results. Major differences are highlighted.
Impact of occlusion-aware parameterization. One of the major contributions that enables Wild2Avatar to achieve human renderings under occlusions is the occlusion-aware scene parameterization. Here we verify the effectiveness of our design by replacing it with an alternative optimization strategy that simply optimizes human NeRF on the visible parts only [8] while maintaining all the objectives except $\mathcal{L}_{occ}$. As shown in the first column of Figure 11, even though Wild2Avatar can still recover the occluded appearances without the proposed parameterization, the rendered results are corrupted with many artifacts due to data scarcity.
Impact of $\mathcal{L}_{occ}$. In section 3.3, the occlusion decoupling loss is proposed to further disentangle the rendering of human and occlusions. We validated the effect of this loss by simply removing it from optimization objectives. As can be seen in the second column of Figure 11, the proposed loss is necessary to fully recover the occluded region of the human.
Impact of $\mathcal{L}_{comp}$. As noted in [8], the completeness of the 3D human geometry needs to be explicitly enforced during training. When removing the completeness loss, the 2D renderings can easily degenerate. We also observed that the proposed loss serves as a regularizer to force the human geometry to be consistent with the SMPL mesh, which helps prevent the model from rendering incorrect poses (third column in Figure 11).
Impact of optimizing inaccurate SMPL jointly with Wild2Avatar. For our evaluations on both in-the-wild and synthesized videos, SMPL parameters were estimated using an occlusion robust HMR model [63]. As shown in Figure 6 and Figure 7, Wild2Avatar demonstrates strong generalization to inaccurate SMPL parameters. To further validate this, we conducted additional experiments on OcMotion videos, jointly optimizing predicted SMPL parameters and Wild2Avatar. As highlighted by the red arrows in Figure 13, this joint optimization progressively refines the SMPL poses, ultimately achieving reconstruction quality comparable to that obtained using ground-truth SMPL parameters.
@@ 4.7 Visualizations of Scene Decomposition
4.8 Generative Ability
While existing methods to render occluded humans like OccNeRF [8] assume that every part of the human body needs to be visible at some point of the sequence, Wild2Avatar is able to hallucinate reasonable appearances for unobserved human parts. This is due to the design of the three-stage scene parameterization that enables standard photometric fitting on both visible and invisible human parts without forcing the model to overfit on the visible parts. We conducted supportive experiments to showcase Wild2Avatar’s generative ability by gradually removing frames that show the back of the human. As shown in Figure 12, Wild2Avatar still renders a complete and reasonable human geometry even without a clear view of the human’s back.
Figure 12: Study of generative ability. Qualitative comparison with Vid2Avatar [1] on amount of training images that show the full back of the human.
Figure 13: Comparison between joint training with predicted SMPL parameters (left) and GT parameters (right). Inaccurate limb rotations in the predicted SMPL can be progressively corrected through joint optimization.
@@ 5 Discussion and Conclusion
Discussion. Training human NeRFs to fit input frames while handling occlusions is challenging. Without any semantic priors of the occluders, fully unsupervised occlusion-human decoupling is impractical. Our proposed occlusion decoupling loss $\mathcal{L}_{occ}$ provides a weakly-supervised solution using precomputed binary human segmentation masks. Both Vid2Avatar and Wild2Avatar rely on accurate pose estimates, which can lead to poor or incomplete rendering if human poses are initialized inaccurately. Moreover, modeling occlusions increases inference time, slowing down the optimization process. While these challenges persist, human pose estimation under occlusion is an actively evolving field, with new occlusion-robust methods such as PARE [64], 3DNBF [65], DPMesh [63], GLAMR [66], and SLAHMR [60] continuing to improve robustness and temporal consistency. Our method builds on these advancements and offers a practical and effective solution for high-quality reconstruction of human geometry and appearance under occlusions. Future work may incorporate additional pose correction steps [60] and efficient NeRF variants [22, 67, 30] or Gaussian splatting [30] for faster training.
@@ Acknowledgment.
This work was partially funded by the NIH Grant R01AG089169 and P41EB027060, Panasonic Holdings Corporation, the Gordon and Betty Moore Foundation, the Jaswa Innovator Award, Stanford HAI, and Stanford Wu Tsai Human Performance Alliance.
6 Biography
Tiange Xiang is a CS Ph.D. candidate at Stanford University. He is affiliated with Stanford AI Lab & Stanford Vision and Learning Lab. He is advised by Prof. Fei-Fei Li, co-advised by Prof. Scott Delp and Prof. Ehsan Adeli. His research focuses on 3D computer vision & AI for healthcare, which is partially supported by Stanford Human-centered AI fellowship. .
Adam Sun is a senior and 1st-year master’s student at Stanford University. He is interested in 3D computer vision for scene understanding and medical applications. Adam is a student researcher at Stanford Vision and Learning Lab (SVL), where he is advised by Professors Fei-Fei Li and Ehsan Adeli at the Partnership in AI-Assisted Care.
Dr. Scott Delp is the James H. Clark Professor at Stanford University, where he holds appointments in Bioengineering, Mechanical Engineering, and Orthopaedic Surgery. He is the founding chair of Stanford’s Bioengineering Department and directs major initiatives including the Wu Tsai Human Performance Alliance, the Restore Center, and the Mobilize Center. A pioneer in movement science and digital health, he developed widely-used tools like OpenSim and has authored over 325 research papers.
Dr. Kazuki Kozuka is a manager at Panasonic Corporation. He received his Ph.D degrees from Nagoya Institite of Technology in 2009. In 2009, he joined Advanced Technology Laboratories, Panasonic Corporation, Japan, where he was involved in research on neuroscience, image recognition. He was also a visiting scholar in Stanford University between 2016 and 2019. His research interests lie in visual understanding through machine learning, mainly for computer vision and natural language processing.
Dr. Fei-Fei Li is the inaugural Sequoia Professor in the Computer Science Department at Stanford University, and Co-Director of Stanford’s Human-Centered AI Institute. She served as the Director of Stanford’s AI Lab from 2013 to 2018. And during her sabbatical from Stanford from January 2017 to September 2018, Dr. Li was Vice President at Google and served as Chief Scientist of AI/ML at Google Cloud. Since then she has served as a Board member or advisor in various public or private companies.
Dr. Ehsan Adeli is an assistant professor at Stanford University School of Medicine, Department of Psychiatry and Behavioral Sciences, where he directs the Stanford Translational Artificial Intelligence (STAI) in Medicine and Mental Health Lab. He is also affiliated with the computer science department, Stanford Vision and Learning (SVL) lab. With a Ph.D. in computer science & artificial intelligence and postgraduate training in biomedical imaging & computational neuroscience, Ehsan is applying his expertise to solve critical problems in healthcare and neuroscience. To this end, his team focuses on developing innovative methods to enhance our understanding of digital humans within the 3D world.